# Sharp thresholds for spanning regular subgraphs

### v2

@@ 4. Fragmentation: arbitrary families
Let $\delta_{n}=o(1)$ be a slowly decreasing function. Let $d\geq 3$ be a constant,
$p=p(n)=\Theta(n^{-2/d}),\quad\quad m=m(n)=pN,\quad\quad f=f(n)=O(n).$
@@ 2.1. If $c<\varepsilon^{\prime}x$ and $x<\varepsilon^{\prime}n$, then
It is well known that increasing properties that hold whp in the uniform model hold whp in the respective binomial model as well, and vice versa (see, e.g., [17, Corollary 1.16]). Let
@@
By [17, Corollary 1.16], whp there exists a multiset $\mathcal{F}^{(\lceil 1/\delta\rceil)}_{n}=\mathcal{F}^{(\lceil 1/\delta\rceil)}_{n}(\mathbf{G})$ of graphs of size $\lfloor\ln n\rfloor$ comprising a single subgraph $H$ of almost every $F\in\mathcal{F}_{n}$ so that $H\cup\mathbf{G}$ contains some $F^{\prime}\in\mathcal{F}_{n}$, since the existence of such a multiset is an increasing property.
@@ 5.2.1. Fragmentation: two iterations
Now, let us bound the number of ways to choose a subgraph $J\subset F$ such that $J\in\mathcal{J}_{\ell,x,c,c^{\prime}}$. The number of ways to choose $c$ disjoint edges in $F$ is at most
@@
We conclude that, for some constants $A_{1},A_{2}>0$,
@@
$\frac{\varphi_{1}\varphi_{2}(n-x+c^{\prime})!}{n!}$ $\leq 4c^{-1/2}\frac{n}{n-2c+1}\cdot\frac{\left(\frac{ed^{2}n^{2}}{2c}\right)^{c}}{n^{2\ell/d+(2-2/d)c+c^{\prime}+2(\sigma-(d-\Delta/2)c^{\prime})/d}}\cdot e^{\frac{(x-c^{\prime})^{2}}{n}}.$
The latter implies
$\frac{\varphi_{1}\varphi_{2}(n-x+c^{\prime})!}{n!}e^{-\frac{(nd/2)^{2}}{m}}\left(\frac{N}{m}\right)^{c}$ $\leq\frac{4n}{n-2c+1}\cdot\frac{c^{-1/2}\left(\frac{ed^{2}n^{2}n^{2/d}}{4cm}\right)^{c}}{n^{2\ell/d+c^{\prime}+2(\sigma-(d-\Delta/2)c^{\prime})/d}}\cdot e^{-\frac{(nd/2)^{2}}{m}+\frac{(x-c^{\prime})^{2}}{n}}$
@@
1. $\ell\leq Cn^{2/d}$. We get
$p(\ell,x,c,c^{\prime})e^{-\frac{(nd/2)^{2}}{m}}\left(\frac{N}{m}\right)^{\ell}\left(1+\frac{3dn}{2m}\right)^{\ell}=O\left(\frac{{n-2c\choose c^{\prime}}{x-2c\choose c^{\prime}}e^{A_{1}c^{\prime}+A_{2}\sigma}{x-2c\choose o}^{2}\left(\frac{N}{m}\right)^{\ell-c}}{\sqrt{c}\cdot n^{2(\ell-c)/d+c^{\prime}+2(\sigma-(d-\Delta/2)c^{\prime})/d}}\right),$ (24)
@@
1.1. If $\sigma<\varepsilon^{\prime}(x-2c)$, then, since $c^{\prime}\leq\frac{2}{2d-\Delta}\sigma$ and $\ell\geq x-c-c^{\prime}$, we get
@@
$<\left(\frac{e^{1+2/d}}{c^{\prime}}\right)^{c^{\prime}}e^{-\frac{x-2c}{d}}\cdot c^{-1/2}.$
@@
$\leq e^{\tau}\sum_{t}e^{t-(\tau+t)t/\tau+(t/\tau)^{2}(\tau+t)/2}\leq e^{\tau}\sum_{t}e^{-(1-\varepsilon^{\prime})t^{2}/(2\tau)}<3\sqrt{\tau}\cdot e^{\tau}.$
@@
$\left(\frac{ed^{2}n^{2}}{4cm}\right)^{c}e^{-\frac{(nd/2)^{2}}{m}}=e^{-\Omega(n^{2/d})}.$
Summing up, we get that, for every fixed $c^{\prime}$ and $x-2c$,
$\sum_{c}p(\ell,x,c,c^{\prime})e^{-\frac{(nd/2)^{2}}{m}}\left(\frac{N}{m}\right)^{\ell}\left(1+\frac{3dn}{2m}\right)^{\ell}=O\left(\left(\frac{e^{1+2/d}}{c^{\prime}}\right)^{c^{\prime}}\cdot e^{-(x-2c)/d}\right)+e^{-\Omega(n^{2/d})}.$
Since for a fixed $c$ and fixed $x-2c$, the value of $\ell$ ranges between $c+(x-2c-c^{\prime})$ and $c+(x-2c-c^{\prime})\frac{d}{2}$, we get
@@ 1.2. If $\sigma\geq\varepsilon^{\prime}(x-2c)$ and $c^{\prime}\leq\frac{1}{2d-\Delta}\sigma$, then
Let $\mathbf{G}\sim G(n,p)$, where $p=(1+2\varepsilon)(e/n)^{2/d}$. By [17, Corollary 1.16], whp there exists a multiset $\mathcal{F}^{*}_{n}=\mathcal{F}^{*}_{n}(\mathbf{G})$ of graphs of size $\lfloor\ln n\rfloor$ comprising a single subgraph $H$ of almost every $F\in\mathcal{F}_{n}$ so that $H\cup\mathbf{G}$ contains some $F^{\prime}\in\mathcal{F}_{n}$ and $H$ has at most $\sqrt{\ln n}$ non-isolated edges.
@@
As noted in Section 1.4 (and evident from the current proof), we have $\mathrm{Var}X_{F}=O(({\sf E}X_{F})^{2})$. Consequently, Theorem 5.4 follows from Paley–Zygmund inequality and [13, Theorem 2.3]. However, this holds only when there are very few subgraphs whose edge boundary is smaller than $2d$. In particular, for $d=3$, there are at least linearly many subgraphs that do not satisfy this condition. Specifically, any pair of incident edges forms a subgraph on three vertices with an edge boundary of at most five. This observation prevents the direct application of both Riordan’s second-moment argument [31] and the fragmentation method in its current form, necessitating new ideas.
@@
For every $\varepsilon>0$, whp $\mathbf{G}\sim G\left(n,(1+\varepsilon)\sqrt{e/n}\right)$ contains the second power of a Hamilton cycle.
@@
Everywhere in this section $F$ is a second power of a cycle on $[n]$ and $\mathcal{F}_{n}$ is the set of all isomorphic copies of $F$ on $[n]$. Let us recall that we call a subgraph $H\subset F\in\mathcal{F}_{n}$ closed if it is the second power of a path.
The proof of Theorem 6.1 is organised as follows. In Section 6.1, we define a subfamily of $\mathcal{F}_{n}$ that is convenient for improving fragments: we fix $\omega(\sqrt{n}/\ln n)$ diamonds (4-vertex graphs obtained by removing a single edge from a 4-clique) and include into the family all $F\in\mathcal{F}_{n}$ that contain these diamonds at equal distances. Then we prove that typical graphs from this subfamily have a fragment of size $O(\sqrt{n})$. In Section 6.2, we improve fragments by moving large closed subgraphs ‘inside’ the diamonds and then prove that there exists a way to assign closed subgraphs to the diamonds so that the multiset of improved fragments have “small spreads” — see Claim 6.7. We state and prove this reconstruction claim in Section 6.4. One of the main ingredients of this proof is the novel strategy of selecting a fragment, which is presented in Section 6.3. Due to Claim 6.7, in the next fragmentation round in Section 6.5 we get fragments of size $O(n^{0.3})$. In Section 6.6, we show that the usual fragmentation technique, equipped with the reconstruction claim, allows to reduce the size of fragments to $o(\log n)$. Finally, in Section 6.7, we complete the proof: we show that sprinkling $\varepsilon p{n\choose 2}$ extra edges is enough to cover at least one remaining fragment. In this last step, the reconstruction claim is one of the main ingredients, as well. Crucially, its second part — the inequality (39) — is used to cover small fragments in this last step of the proof, and the proof of (39) essentially relies on the distribution of a random fragment, presented in Section 6.3.
@@ 6.1. Day 0: getting fragments of size $O(\sqrt{n})$
In what follows, for the sake of convenience, $\mathcal{F}_{n}$ is the family of all rooted oriented second powers of a cycle on $[n]$, that is, each cycle $F\in\mathcal{F}_{n}$ has a specified vertex — its root — and is oriented, starting from the root, in one of the two directions. Clearly, $|\mathcal{F}_{n}|=n!$ and each $F\in\mathcal{F}_{n}$ induces a linear order on $[n]$; we denote the respective permutation acting on $[n]$ by $\pi_{F}$.
@@
  • the order of the diamonds as well as the order of vertices in the diamonds is aligned with $\pi_{F}$ and the first vertex of $D_{1}$ coincides with the root of $F$,
@@
Fix $F\in\mathcal{F}_{n}(\overrightarrow{D})$ and $W\in{{[n]\choose 2}\choose m}$. An $(F,W)$-fragment is a subgraph $H$ of some $F^{\prime}\in\mathcal{F}_{n}(\overrightarrow{D})$, such that $F^{\prime}\subset F\cup W$ and $H=F\cap F^{\prime}$. We will show that still whp there exists an $(F,\mathbf{W})$-fragment $H$ of size $O(\sqrt{n})$. In particular, this fragment has to contain all diamonds $D_{1},\ldots,D_{\chi}$. This will allow us to refine the fragment $H$ by spreading evenly the edges of large closed subgraphs in $H$ among the diamonds and, thus, to get rid of all the closed subgraphs of size $\Omega(\ln n)$.
@@
We claim that whp $|\mathcal{M}(F,\mathbf{W})|=\Omega(M)$ and that typically most $(F,\mathbf{W})$-fragments have size $O(\sqrt{n})$. Although the conclusion of Lemma 4.1 is sufficient for our goals, below we state its stronger version which holds for symmetric families $\mathcal{F}_{n}$ (i.e. when all graphs in the family are isomorphic). In this version, we replace a random element of $\mathcal{F}^{*}_{n}(\overrightarrow{D})$ with any fixed element. Since the proofs are almost identical, we postpone them to Appendix B.
@@
$\sum_{\ell\geq\ell_{0}+1}\Pi_{\ell}\cdot e^{-\frac{4n^{2}}{m}}\left((1+o(1))\frac{N}{m}\right)^{\ell}=o\left(\frac{1}{n}\right).$ (34)
@@
[…]
@@
Let $F\in\mathcal{F}^{*}_{n}(\overrightarrow{D})$ and let $\mathbf{W}\subset{[n]\choose 2}\setminus(D_{1}\cup\ldots\cup D_{\chi})$ be a uniformly random $m$-subset. Then whp $(F,\mathbf{W})$ is not $\ell_{0}$-bad.

### v3

@@ 4. Fragmentation: arbitrary families
Let $\delta_{n}=o(1)$ be a slowly decreasing function. Let
$p=p(n)=\Omega(n^{-2/3}),\quad\quad m=m(n)=pN,\quad\quad f=f(n)=O(n).$
@@ 2.1. If $c<\varepsilon^{\prime}x$ and $x<\varepsilon^{\prime}n$, then
It is well known that increasing properties that hold whp in the uniform model hold whp in the respective binomial model as well, and vice versa (see, e.g., [18, Corollary 1.16]). Let
@@
By [18, Corollary 1.16], whp there exists a multiset $\mathcal{F}^{(\lceil 1/\delta\rceil)}_{n}=\mathcal{F}^{(\lceil 1/\delta\rceil)}_{n}(\mathbf{G})$ of graphs of size $\lfloor\ln n\rfloor$ comprising a single subgraph $H$ of almost every $F\in\mathcal{F}_{n}$ so that $H\cup\mathbf{G}$ contains some $F^{\prime}\in\mathcal{F}_{n}$, since the existence of such a multiset is an increasing property.
@@ 5.2.1. Fragmentation: two iterations
Now, let us bound the number of ways to choose a subgraph $J\subset F$ such that $J\in\mathcal{J}_{\ell,x,c,c^{\prime}}$. The number of ways to choose $c>0$ disjoint edges in $F$ is at most
@@
Letting $\varphi_{2}:=1$ when $c=0$, we conclude that, for all $c\geq 0$ and for some constants $A_{1},A_{2}>0$,
@@
[…]
In particular, when $c>0$, we get
$\frac{\varphi_{1}\varphi_{2}(n-x+c^{\prime})!}{n!}e^{-\frac{(nd/2)^{2}}{m}}\left(\frac{N}{m}\right)^{c}$ $\leq\frac{4n}{n-2c+1}\cdot\frac{c^{-1/2}\left(\frac{ed^{2}n^{2}n^{2/d}}{4cm}\right)^{c}}{n^{2\ell/d+c^{\prime}+2(\sigma-(d-\Delta/2)c^{\prime})/d}}\cdot e^{-\frac{(nd/2)^{2}}{m}+\frac{(x-c^{\prime})^{2}}{n}}$ (23)
@@
1. $\ell\leq Cn^{2/d}$. Since $d\geq 4$, the equalities $n\ell=O(m)$ and $(x-c^{\prime})^{2}=O(n)$ hold. In particular, $\left(1+\frac{3dn}{2m}\right)^{\ell}e^{(x-c^{\prime})^{2}/n}=O(1)$. Combining (19), (22), and (24), we get
$p(\ell,x,c,c^{\prime})e^{-\frac{(nd/2)^{2}}{m}}\left(\frac{N}{m}\right)^{\ell}\left(1+\frac{3dn}{2m}\right)^{\ell}\\ =O\left(\frac{{n-2c\choose c^{\prime}}{x-2c\choose c^{\prime}}e^{A_{1}c^{\prime}+A_{2}\sigma}}{c^{\prime}+2(\sigma-(d-\Delta/2)c^{\prime})/d}\left(\frac{{x-2c\choose o}^{2}\left(\frac{N}{m}\right)^{\ell-c}}{\sqrt{c}\cdot n^{2(\ell-c)/d}}\cdot\mathbbm{1}_{c>0}+\frac{{x\choose o}^{2}\left(\frac{N}{m}\right)^{\ell}}{n^{2\ell/d}}e^{-\frac{(nd/2)^{2}}{m}}\cdot\mathbbm{1}_{c=0}\right)\right),$ (26)
@@
1.1. If $\sigma<\varepsilon^{\prime}(x-2c)$, then, since $c^{\prime}\leq\frac{2}{2d-\Delta}\sigma$ and $\ell\geq x-c-c^{\prime}$, we get, for $c>0$,
@@
$<\left(\frac{e^{1+2/d}}{c^{\prime}}\right)^{c^{\prime}}e^{-\frac{x-2c}{d}}\cdot c^{-1/2}.$ (27)
@@
$\leq e^{\tau}\sum_{t}e^{t-(\tau+t)t/\tau+(t/\tau)^{2}(\tau+t)/2}\leq e^{\tau}\sum_{t}e^{-(1-\varepsilon^{\prime})t^{2}/(2\tau)}<3\sqrt{\tau}\cdot e^{\tau}.$ (28)
@@
$\left(\frac{ed^{2}n^{2}}{4cm}\right)^{c}e^{-\frac{(nd/2)^{2}}{m}}=e^{-\Omega(n^{2/d})}.$ (29)
We are now ready to estimate the contribution of the considered values of parameters $\ell,x,c,c^{\prime}$ to the left-hand side of (25). Let us again recall the inequality (19). In order to refine the bound (27), we will use the bound (23) instead of (24). Recalling that $c^{\prime}\leq\frac{2}{2d-\Delta}\sigma$ and arguing in a similar way as in (27), we get that any summand in the left-hand side of (25) that corresponds to $c>0$ is at most
$\frac{\varphi_{1}\varphi_{2}{n-2c\choose c^{\prime}}{x-2c\choose c^{\prime}}e^{A_{1}c^{\prime}+A_{2}\sigma}(n-x+c^{\prime})!}{n!}{x-2c\choose o}^{2}e^{-\frac{(nd/2)^{2}}{m}}\left(\frac{N}{m}\right)^{\ell}\left(1+\frac{3dn}{2m}\right)^{\ell}\\ =O\left(\frac{c^{-1/2}e^{A_{1}c^{\prime}+A_{2}\sigma}}{n^{2(\ell-c)/d+c^{\prime}}}\cdot\left(\frac{e(nd/2)^{2}}{cm}\right)^{c}e^{-\frac{(nd/2)^{2}}{m}}{n-2c\choose c^{\prime}}{x-2c\choose c^{\prime}}{x-2c\choose o}^{2}\left(\frac{N}{m}\right)^{\ell-c}\right)\\ =O\left(\left(\frac{e^{1+2/d}}{c^{\prime}}\right)^{c^{\prime}}e^{-\frac{x-2c}{d}}\cdot c^{-1/2}\left(\frac{e(nd/2)^{2}}{cm}\right)^{c}\cdot e^{-\frac{(nd/2)^{2}}{m}}\right).$
Then, for every fixed $c^{\prime}$ and $x-2c$, from (28) we derive
$\sum_{c\in[(1-\varepsilon^{\prime})\tau,(1+\varepsilon^{\prime})\tau]}p(\ell,x,c,c^{\prime})e^{-\frac{(nd/2)^{2}}{m}}\left(\frac{N}{m}\right)^{\ell}\left(1+\frac{3dn}{2m}\right)^{\ell}=O\left(\left(\frac{e^{1+2/d}}{c^{\prime}}\right)^{c^{\prime}}\cdot e^{-(x-2c)/d}\right).$
Due to (29), we get that $c\notin[(1-\varepsilon^{\prime})\tau,(1+\varepsilon^{\prime})\tau]$ contribute $e^{-\Omega(n^{2/d})}$ to the left-hand side of (25), implying
$\sum_{c\geq 0}p(\ell,x,c,c^{\prime})e^{-\frac{(nd/2)^{2}}{m}}\left(\frac{N}{m}\right)^{\ell}\left(1+\frac{3dn}{2m}\right)^{\ell}=O\left(\left(\frac{e^{1+2/d}}{c^{\prime}}\right)^{c^{\prime}}\cdot e^{-(x-2c)/d}\right)+e^{-\Omega(n^{2/d})}.$
Since for fixed $c$ and fixed $x-2c$, the value of $\ell$ ranges between $c+(x-2c-c^{\prime})$ and $c+(x-2c-c^{\prime})\frac{d}{2}$, we get
@@ 1.2. If $\sigma\geq\varepsilon^{\prime}(x-2c)$ and $c^{\prime}\leq\frac{1}{2d-\Delta}\sigma$, then
Let $\mathbf{G}\sim G(n,p)$, where $p=(1+2\varepsilon)(e/n)^{2/d}$. By [18, Corollary 1.16], whp there exists a multiset $\mathcal{F}^{*}_{n}=\mathcal{F}^{*}_{n}(\mathbf{G})$ of graphs of size $\lfloor\ln n\rfloor$ comprising a single subgraph $H$ of almost every $F\in\mathcal{F}_{n}$ so that $H\cup\mathbf{G}$ contains some $F^{\prime}\in\mathcal{F}_{n}$ and $H$ has at most $\sqrt{\ln n}$ non-isolated edges.
@@
As noted in Section 1.5 (and evident from the current proof), we have $\mathrm{Var}X_{F}=O(({\sf E}X_{F})^{2})$. Consequently, Theorem 5.4 follows from Paley–Zygmund inequality and [14, Theorem 2.3]. However, this holds only when there are very few subgraphs whose edge boundary is smaller than $2d$. In particular, for $d=3$, there are at least linearly many subgraphs that do not satisfy this condition. Specifically, any pair of incident edges forms a subgraph on three vertices with an edge boundary of at most five. This observation prevents the direct application of both Riordan’s second moment argument [33] and the fragmentation method in its current form, necessitating new ideas.
@@
For every $\varepsilon>0$, whp $\mathbf{G}\sim G\left(n,(1+\varepsilon)\sqrt{e/n}\right)$ contains the square of a Hamilton cycle.
@@
Everywhere in this section $F$ is the square of a cycle on $[n]$ and $\mathcal{F}_{n}$ is the set of all isomorphic copies of $F$ on $[n]$. Let us recall that we call a subgraph $H\subset F\in\mathcal{F}_{n}$ closed if it is the square of a path.
The proof of Theorem 6.1 is organised as follows. In Section 6.1, we define a subfamily of $\mathcal{F}_{n}$ whose typical fragments contain sufficiently many disjoint powers of paths, and thus is amenable to improving fragments: we fix $\omega(\sqrt{n}/\ln n)$ diamonds (4-vertex graphs obtained by removing a single edge from a 4-clique) and include into the family all $F\in\mathcal{F}_{n}$ that contain these diamonds at equal distances. Then we prove that typical [^7] In the measure induced by $\mathbf{G}$. graphs from this subfamily have fragments of size $O(\sqrt{n})$. In Section 6.2, we improve fragments, chosen from every typical $F\in\mathcal{F}_{n}$ according to certain distribution that is defined later in Section 6.3, by moving large closed subgraphs ‘inside’ the diamonds. In the same section, we describe a probabilistic approach to assign closed subgraphs to the diamonds. It is designed in a way so that the obtained multiset of improved fragments have “small spreads” — see Claim 6.7. We state and prove this reconstruction claim in Section 6.4. One of the main ingredients of this proof is the novel strategy of selecting a fragment, which is presented in Section 6.3.
Due to Claim 6.7, in the second fragmentation round presented in Section 6.5 we get fragments of size $O(n^{0.3})$. In Section 6.6, we show that the usual fragmentation technique that relies on typical fragments, equipped with the reconstruction claim, allows to reduce the size of fragments to $o(\log n)$ in $O(\log\log n)$ rounds, where in each round we expose $O(pn^{2}/\log\log n)$ edges. Finally, in Section 6.7, we complete the proof: we show that sprinkling $\varepsilon p{n\choose 2}$ extra edges is enough to cover at least one remaining fragment. In this last step, the reconstruction claim is one of the main ingredients, as well. Crucially, the second part of the reconstruction claim — the inequality (44) — is used to cover small fragments in this last step of the proof, and the proof of (44) essentially relies on the distribution of a random fragment, presented in Section 6.3.
@@ 6.1. Day 0: getting fragments of size $O(\sqrt{n})$
In what follows, for the sake of convenience, $\mathcal{F}_{n}$ is the family of all rooted oriented squares of cycles on $[n]$, that is, each cycle $F\in\mathcal{F}_{n}$ has a specified vertex — its root — and is oriented, starting from the root, in one of the two directions. Clearly, $|\mathcal{F}_{n}|=n!$ and each $F\in\mathcal{F}_{n}$ induces a linear order on $[n]$; we denote the respective permutation acting on $[n]$ by $\pi_{F}$.
@@
  • the order of the diamonds as well as the order of vertices in the diamonds respects $\pi_{F}$ and the first vertex of $D_{1}$ coincides with the root of $F$,
@@
Fix $F\in\mathcal{F}_{n}(\overrightarrow{D})$ and $W\in{{[n]\choose 2}\choose m}$. An $(F,W)$-fragment is a subgraph $H$ of some $F^{\prime}\in\mathcal{F}_{n}(\overrightarrow{D})$, such that $F^{\prime}\subset F\cup W$ and $H=F\cap F^{\prime}$. We will show that whp there exists an $(F,\mathbf{W})$-fragment $H$ of size $O(\sqrt{n})$. In particular, this fragment has to contain all diamonds $D_{1},\ldots,D_{\chi}$. This will allow us to refine the fragment $H$ by spreading evenly the edges of large closed subgraphs in $H$ among the diamonds and, thus, to get rid of all the closed subgraphs of size $\Omega(\ln n)$.
@@
We claim that $|\mathcal{M}(F,\mathbf{W})|=\Omega_{p}(M)$ and that typically most $(F,\mathbf{W})$-fragments have size $O(\sqrt{n})$. Although the conclusion of Lemma 4.1 is sufficient for our goals, below we state its stronger version which holds for symmetric families $\mathcal{F}_{n}$ (i.e. when all graphs in the family are isomorphic). In this version, we replace a random element of $\mathcal{F}^{*}_{n}(\overrightarrow{D})$ with any fixed element. Since the proofs are almost identical, we postpone them to Appendix B.
@@
$\sum_{\ell\geq\ell_{0}+1}\Pi_{\ell}\cdot e^{-\frac{4n^{2}}{m}}\left((1+o(1))\frac{N}{m}\right)^{\ell}=\exp(-\Omega(\sqrt{n})).$ (39)
@@
For $\ell\in\mathbb{N}$, let $\mathcal{M}_{\ell}(F,W)$ be the set of all $F^{\prime}\in\mathcal{M}(F,W)$ such that $|F^{\prime}\cap F|>\ell$. Let us say that $(F,W)$ is $\ell$-bad, if $|\mathcal{M}_{\ell}(F,W)|>\frac{1}{\sqrt{n}}|\mathcal{M}(F,W)|$. In a similar way as in the proof of Lemma 4.1, we derive from Claim 39 that $(F,\mathbf{W})$ is not $\ell_{0}$-bad whp.
@@
Let $F\in\mathcal{F}^{*}_{n}(\overrightarrow{D})$ and let $\mathbf{W}\subset{[n]\choose 2}\setminus(D_{1}\cup\ldots\cup D_{\chi})$ be a uniformly random $m$-subset. Then $(F,\mathbf{W})$ is not $\ell_{0}$-bad with probability at least $1-\exp(-\Omega(\sqrt{n}))$.